Find the expected count and the contribution to the chi-square statistic for the (Group 2, Yes) cell in the two-way table below.

Yes No Total
Group 1 59 42 101
Group 2 136 64 200
Group 3 67 28 95
Total 262 134 396



Round your answer for the excepted count to one decimal place, and your answer for the contribution to the chi-square statistic to three decimal places.


Expected count = Enter your answer; Expected count

contribution to the chi-square statistic = Enter your answer; contribution to the chi-square statistic
Answer #1

Answer

Expected value = (product of column total and row total)/(grand total)

here row total = 200 and column total = 262

grand total = 396

So, E[x] for group 2 yes is

E[x] =(200*262)/396

= 132.3

Contribution to chi square test statistic

Contribution = (O-E)^2/E

where O is observed value = 136 and E is expected value = 132.3

So, contribution = (136-132.3)^2/132.3

= 0.103